OpenStudy (anonymous):

The length of a rectangle is 6cm longer than its width. If the perimeter of the rectangle is 36cm , find its area.

OpenStudy (anonymous):



OpenStudy (anonymous):

A = w * (w + 6) = 6 * 12 = .....

OpenStudy (nurali):

Let x = width of the rectangle x+6 = length of the rectangle Perimeter = 2(W) + 2(L) = 36 2(x) + 2(x+6) = 36 2x + 2x + 12 = 36 4x + 12 = 36 4x = 24 x= 6 cm Width x+6 = 12 cm Length Area = L*W = 12 cm * 6 cm = ----------sq cm.